# Finite Population Identification and

### v2

@@ 2.2 Identified sets for ATE
To simplify the exposition, from here on we assume that all units in the population are observed, and focus solely on the uncertainty due to missing potential outcomes. In section 7.1 we show that the extension to incorporate sampling is straightforward. We also focus on the identification of the average treatment effect for brevity. We briefly discuss other parameters in section 7.3.
It is well known that, because averages are influenced by the values of outlier observations, bounds on ATE are usually infinite without some kind of restriction on the magnitude of outliers. Hence we maintain the following assumption throughout the paper.
@@
There are known values $-\infty<y_{\text{min}}<y_{\text{max}}<\infty$ such that $Y_{i}(x)\in[y_{\text{min}},y_{\text{max}}]$ for all $i\in\mathcal{I}$, for each $x\in\{0,1\}$.
@@
A1 alone has identifying power for ATE, via a finite population version of the Manski (1990) bounds. This result is a special case of a more general result that uses the following exogeneity assumption.
@@
Suppose A1 and A2 hold, and $\mathbf{P}^{\text{data}}$ is known. Then, for each $x\in\{0,1\}$, the identified set for $\overline{\mathbf{Y}(x)}$ is $[\text{LB}_{K}(x),\text{UB}_{K}(x)]$ where
@@ 2.3 Randomization and identification in finite populations
[…]
$\frac{1}{N_{1}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=1)=\frac{1}{N_{0}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=0)$
for $x\in\{0,1\}$. In fact, for many values of the vectors $\mathbf{Y}(x)=(Y_{1}(x),\ldots,Y_{N}(x))$, it is impossible for exact balance to hold regardless of the values of realized treatment $\mathbf{X}=(X_{1},\ldots,X_{N})$, a point noted by Greenland and Robins (1986, page 415). So statistical independence is not an appropriate formalization of random assignment in finite populations. Instead, we make the following assumption.
@@
The size of the treatment and control groups, $N_{1}$ and $N_{0}$, are fixed a priori. $\mathbf{X}=(X_{1},\ldots,X_{N})$ is a single realization from the known probability distribution $\mathbb{P}_{\text{design}}$ on $\{0,1\}^{N}$ defined by
@@
As has long been recognized, randomization in finite populations does not guarantee any level of ex post balance (e.g., Greenland 1990). The identified set, by definition, must contain the true parameter whenever the model is true. Consequently, for any finite population, randomization has no identifying power. The following theorem formalizes this result.
@@
Put differently, after treatment has been assigned, we cannot rule out the possibility that potential outcomes are substantially imbalanced across the treatment and control groups, even if it was unlikely a priori. Thus, from an identification perspective, randomization does not shrink the domain-of-consensus bounds in finite populations. Nonetheless, in the next section we’ll show that theorem 2 provides an overly pessimistic view of the value of randomization. There we reinterpret randomization as a procedure that affects our beliefs about ex post balance. Specifically, we will use randomization to assess the plausibility of a specific choice of the sensitivity parameter $K$. This will let us use the identification result in theorem 1 to perform a sensitivity analysis motivated by random assignment of treatment.
@@ 3.2 Interpreting $\Theta_{I}(K(\alpha))$
Suppose A1 holds. Then for any realization $\mathbf{X}$, $\underline{p}$ is monotonic, $\underline{p}$ is right continuous, $\lim_{K\rightarrow 0}\;\underline{p}(K)=0$, and $\lim_{K\rightarrow\infty}\;\underline{p}(K)=1$.
$K(\alpha)$ is therefore the smallest value of $K$ such that $K^{\text{true,max}}\leq K$ occurs with probability at least $1-\alpha$, according to the “robust” prior $\underline{p}$, which implies the same would be true for any other prior consistent with our baseline assumptions about the potential outcomes. So $[0,K(\alpha)]$ is a Bayesian credible set for $K^{\text{true,max}}$ with respect to $\underline{p}$; it is a set which contains $K^{\text{true,max}}$ with probability about $1-\alpha$ (the true probability is slightly higher than $1-\alpha$ since $\underline{p}$ is not continuous). And, again, this would also be true for any other prior consistent with our baseline assumptions. Consequently, $\Theta_{I}(K(\alpha))$ can be interpreted as a robust $(1-\alpha)$-credible set for the identified set for ATE. Since the identified set must contain the true ATE when the model is not falsified, this implies that $\Theta_{I}(K(\alpha))$ is also a robust $(1-\alpha)$-credible set for ATE itself.
@@
Keep in mind that $\mathbf{X}^{\text{new}}$ is the only random vector here. Theorem 3 shows that $\Theta_{I}(K(\alpha))$ is a $100(1-\alpha)$design-based confidence set. That is, across repeated random assignments of treatment, it will contain the true parameter with probability at least $1-\alpha$. In particular, this result holds for any fixed, finite population size $N$ that satisfies A1.
@@ 3.3 The value of randomization
[…]
@@ 3.4 Measuring the strength of evidence: $\underline{p}(K^{\text{bp}})$ as an alternative to $p$-values
$p$-values are commonly interpreted as quantitative measures of the evidence against the null hypothesis, with small values interpreted as “strong evidence” against the null and larger values interpreted as weaker evidence against it. This is a controversial interpretation (e.g., see the entire 2019 special issue of The American Statistician on “a world beyond $p<0.05$”). In light of this debate about $p$-values, our measure $\underline{p}(K^{\text{bp}})$ can be viewed as an alternative summary statistic that has a well justified interpretation as a quantitative measure of how much evidence the data provides for a specific conclusion. In particular, because our analysis has a Bayesian interpretation, it allows us to make certain probabilistic statements about the true parameter itself. For example, when the naive difference in means estimate is positive, $\underline{p}(K^{\text{bp}})$ is a lower bound on the probability that the true ATE is nonnegative. Here “probability” refers to our randomization-based prior distribution.
@@ 4 Numerical Illustration
Next we use simulated data to illustrate the design-based sensitivity analysis of section 3. We generated a nested sequence of five populations, with $N\in\{10,20,40,100,400\}$. These populations have $[y_{\text{min}},y_{\text{max}}]=[0,1]$ and heterogeneous treatment effects such that $\text{ATE}_{N}$ converges to 0.25 as $N$ gets large. We assigned treatment via A3 such that $N_{1}=N/2$ for all $N$ while ensuring the populations are nested. In this illustration the assignment happens only once. Appendix LABEL:sec:numericalAppendix gives further details on how we generated the data.
[…]

### v3

@@ 2.2 Identified Sets for ATE
To simplify the exposition, we generally assume that all units in the population are observed ($S_{i}=1$ for all $i\in\mathcal{I}$). We show that the extension to incorporate sampling is straightforward in section 7.1. We also focus on the identification of the average treatment effect for brevity. We briefly discuss other parameters in section 7.3.
It is well known that bounds on mean based parameters are usually infinite without some kind of restriction on the values they can take. Hence we maintain the following assumption throughout the paper, which is standard in the literature on partial identification.
@@
There are known values $\infty<y_{\text{min}}<y_{\text{max}}<\infty$ such that $Y_{i}(x)\in[y_{\text{min}},y_{\text{max}}]$ for all $i\in\mathcal{I}$, for each $x\in\{0,1\}$.
@@
A1 alone has identifying power for ATE, via a finite population version of the Manski (1990) bounds. Those bounds are a special case of theorem 1 below, which uses the following assumption.
@@
Suppose A1 and A2 hold, and $\mathbf{P}^{\text{data}}$ is known. Then, for each $x\in\{0,1\}$, the identified set for $\overline{\mathbf{Y}(x)}$ is $\text{LB}_{K}(x),\text{UB}_{K}(x)]$ where
@@ 2.3 Randomization and Identification in Finite Populations
In practice, researchers often motivate assumptions that two groups are in some sense comparable by appealing to random assignment. So consider the following common formalization of random assignment in finite populations.
@@
The size of the treatment and control groups, $N_{1}$ and $N_{0}$, are fixed a priori, with $0<N_{1}<N$. $\mathbf{X}=(X_{1},\ldots,X_{N})$ is a single realization from the known probability distribution $\mathbb{P}_{\text{design}}$ on $\{0,1\}^{N}$ defined by
@@
As has long been recognized, randomization in finite populations does not guarantee any level of ex post balance (e.g., Greenland 1990). The identified set, by definition, must contain the true parameter value whenever the model assumptions are true. Consequently, for any finite population, randomization has no identifying power. The following theorem formalizes this result.
@@
Put differently, after treatment has been assigned, we cannot rule out the possibility that potential outcomes are substantially imbalanced across the treatment and control groups, even if it was unlikely a priori. Thus the assumption that treatment was randomly assigned does not rule out any values of the unknown potential outcomes. Hence it does not shrink the domain-of-consensus bounds in finite populations. Nonetheless, in section 3 we will reinterpret randomization as a procedure that affects our beliefs about ex post balance. Specifically, we will use randomization to assess the plausibility of a specific choice of the sensitivity parameter $K$. This will let us use the identification result in theorem 1 to perform a sensitivity analysis motivated by random assignment of treatment. That analysis will yield uncertainty intervals that have interpretations as identified sets, robust Bayesian credible sets, or uniform frequentist confidence sets.
Exact Balance in Infinite versus Finite Populations
[…]
This qualitative distinction between the finite and infinite setting explains why theorem 2 comes to a different conclusion about identification of ATE from traditional infinite population identification analyses: In the infinite population setting, exact balance is assumed to hold and hence ATE is point identified. But in the finite population setting, exact balance is not guaranteed even under random assignment, and is sometimes even logically impossible, and hence ATE is partially identified. That said, because any level of approximate balance is more likely when $N$ is large, there is a connection between these two results, which we study in section 3.3. Also see section 7.3, where we discuss various measures of balance.
2.4 The Prior Definition of Finite Population Identification
Now that we have explained our approach to identification, we turn to examine the prior definition. First, what is the prior definition? We are not aware of any books or papers that explicitly and formally define identification for finite populations. For example, Borusyak et al. (2025) discuss identification but do not provide a formal definition, and as representative of the rest of the literature, their first formal result is a consistency proof, not an identification theorem. The book by Tillé (2020) is an extensive and detailed survey of sampling from finite populations, but does not discuss identification. The textbook by Ding (2024) discusses causal inference in finite populations in detail, starting from chapter 2, but a definition of identification does not appear until chapter 10, only after shifting focus from randomized experiments to observational data. That definition (10.1 on page 128) states that “a parameter $\theta$ is nonparametrically identifiable if it can be written as a function of the distribution of the observed data without any parametric model assumptions”. Similar definitions appear in other references, going back to the earliest sources (such as Koopmans 1949, Koopmans and Reiersøl 1950, Hurwicz 1950, Rothenberg 1971). Verbal definitions like these are ambiguous about the meaning of “the distribution of the observed data” (Ding 2024) or the “joint distribution of the observations” (Koopmans 1949, page 125). There are different ways to formalize such statements and they lead to substantially different definitions of identification. As we argue below, this issue is particularly important in the finite population setting.
Our definition 1 is one such formalization, where by “data” we mean the matrix $\mathbf{P}^{\text{data}}$ that the researcher actually has at hand. In this case the “distribution” of this data can be formalized as a specific discrete distribution (see appendix LABEL:sec:popDistributions for details). However, based on the informal discussions in the literature, and our personal communication with researchers in this field, “data” is conventionally interpreted as the design distribution of the observed data, rather than the dataset that is actually observed. This interpretation leads to the following definition of point identification for finite populations.
Definition 2 (Conventional).
Let $(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{W})$ be the $N\times(2+d_{W})$ matrix of potential outcomes and covariate values for all units in the population $\mathcal{I}=\{1,\ldots,N\}$. Let $\theta(\cdot)$ be a functional defined on the set of logically possible values of this matrix. Let $\theta^{\text{true}}\coloneqq\theta((\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{W}))$ denote its true value. For each $i\in\mathcal{I}$, let $X_{i}^{\text{new}}\in\{0,1\}$ and $S_{i}^{\text{new}}\in\{0,1\}$ be random variables. Let $\mathbf{X}^{\text{new}}\coloneqq(X_{1}^{\text{new}},\ldots,X_{N}^{\text{new}})$ and $\mathbf{S}^{\text{new}}\coloneqq(S_{1}^{\text{new}},\ldots,S_{N}^{\text{new}})$. Let $\mathbb{P}_{\text{design}}$ denote the joint distribution of $(\mathbf{X}^{\text{new}},\mathbf{S}^{\text{new}})$; it is called the design distribution. Let $Y_{i}^{\text{new}}\coloneqq Y_{i}(X_{i}^{\text{new}})$. Let $\mathcal{S}=\{i\in\mathcal{I}:S_{i}^{\text{new}}=1\}$ denote the random set of sampled indices. Let $\mathcal{D}\coloneqq\{(Y_{i}^{\text{new}},X_{i}^{\text{new}},W_{i})\in\mathbb{R}^{2+d_{W}}:i\in\mathcal{S}\}$. Say $\theta^{\text{true}}$ is point identified if there is a known mapping from the probability distribution of the random set $\mathcal{D}$ to $\theta^{\text{true}}$.
We believe definition 2 accurately formalizes the conventional definition of finite population identification in the prior literature. For example, Heckman and Vytlacil (2007, footnote 6, page 4786) say “identification in small samples requires establishing the sampling distribution of estimators, and adopting bias as the criterion for identifiability” (see our proposition 1 below) and that “this approach is conventional in classical statistics,” although they do not give a formal definition which we could compare against. This definition also resembles the definition of “sampling identification” in Florens and Simoni (2021, definition 2.1 on page 5), although that paper is not about design-based inference or finite populations.
Definition 2 has the following seemingly innocuous and familiar implication.
Proposition 1.
Consider the setting of definition 2. Suppose there is a known function $m$ such that $\theta^{\text{true}}$ is the unique solution to $\mathbb{E}_{\text{design}}[m(\mathcal{D},\theta)]=0$. Then $\theta^{\text{true}}$ is point identified.
This result says that a parameter is point identified if it is the unique solution to a set of design-based moment conditions. A special case arises when there is an $m$ is such that we can write $\theta^{\text{true}}=\mathbb{E}_{\text{design}}[m(\mathcal{D})]$. In this case the function $m(\mathcal{D})$ is usually called a design-unbiased estimator of $\theta^{\text{true}}$. Moment-based reasoning like in proposition 1 is used in the identification analysis of Borusyak et al. (2025, page 88, first paragraph of section 3.1) and Rambachan and Roth (2025, page 13, first displayed equation in second column). Proposition 1, however, implies the following result.
Proposition 2.
Consider the setting of definition 2. Suppose $\mathbb{P}_{\text{design}}(S_{i}^{\text{new}}=1)=1$ for all $i\in\mathcal{I}$. Then, according to definition 2, the following hold.
  1. 1.
    $p_{i}\coloneqq\mathbb{P}_{\text{design}}(X_{i}^{\text{new}}=1)$ is point identified for all $i\in\mathcal{I}$.
  2. 2.
    $Y_{i}(1)$ is point identified for each $i\in\mathcal{I}$ with $p_{i}>0$. $Y_{i}(0)$ is point identified for each $i\in\mathcal{I}$ with $p_{i}<1$.
Proof of proposition 2.
Part 1 follows from $p_{i}=\mathbb{E}_{\text{design}}(X_{i}^{\text{new}})$ and proposition 1. For part 2, define
$\widehat{Y_{i}(1)}\coloneqq\frac{X_{i}^{\text{new}}Y_{i}^{\text{new}}}{p_{i}}\qquad\text{and}\qquad\widehat{Y_{i}(0)}\coloneqq\frac{(1-X_{i}^{\text{new}})Y_{i}^{\text{new}}}{1-p_{i}}.$
Then
$\mathbb{E}_{\text{design}}[\widehat{Y_{i}(1)}]=\frac{1}{p_{i}}\mathbb{E}_{\text{design}}[X_{i}^{\text{new}}Y_{i}(X_{i}^{\text{new}})]\\ =\frac{1}{p_{i}}\mathbb{E}_{\text{design}}[X_{i}^{\text{new}}]Y_{i}(1)\\ =Y_{i}(1).$
Similarly, $\mathbb{E}_{\text{design}}[\widehat{Y_{i}(0)}]=Y_{i}(0)$. Apply proposition 1. ∎
We assumed all units are sampled in proposition 2 for simplicity, but the result can be generalized to allow for sampling. Proposition 2 combines ideas from Imbens (2018, footnote 3) and Chen et al. (2026). Specifically, using a different proof strategy and restricting attention to the binary outcome case, Chen et al. (2026) showed that the proportion of units who benefit from treatment is point identified, according to the conventional definition 2 of identification in finite populations. Imbens (2018) considered a randomized experiment with $N=1$ and $\mathbb{P}(X_{1}^{\text{new}}=1)=0.5$ and showed that there exists a design-unbiased estimator of that units’ unit-level treatment effect, $Y_{1}(1)-Y_{1}(0)$. Our proposition 2 builds on these ideas to show that all potential outcomes are point identified, according to definition 2, as long as the probability of treatment is strictly between zero and one for all units.
Proposition 2 applies to randomized experiments, which typically guarantee $p_{i}\in(0,1)$ by design (e.g., A3 holds), but it also applies to many observational settings as well, because this result does not assume that $p_{i}$ is known for all $i\in\mathcal{I}$. For example, unconfoundedness restricts the treatment assignment probabilities $p_{i}$ to be functions of observed covariates (in which case it is called the propensity score). Proposition 2 implies that such a priori restrictions on treatment probabilities are unnecessary to point identify potential outcomes, according to definition 2; it could even be that the unit-level treatment assignment probabilities $p_{i}=f(Y_{i}(1),Y_{i}(0))$ are functions of potential outcomes themselves; all that matters is that the ex ante probability of treatment is not degenerate on 0 or 1. Put differently, unconfoundedness is unnecessary according to the conventional definition of identification; overlap ($p_{i}\in(0,1)$ for all $i\in\mathcal{I}$) is sufficient for point identification of all unit-level treatment effects. Similarly, the conventional definition implies that all unit-level treatment effects are point identified for compliers in an instrumental variables setting, even if the instrument is endogenous (see appendix LABEL:sec:conventionalAppendix for details).
Essentially all of these conclusions implied by proposition 2 differ from the conclusions of classical infinite population identification theory. The fact that a parameter can be partially identified under infinite population theory but point identified according the conventional definition was first pointed out by Chen et al. (2026), who focused on the proportion of units who benefit from treatment with a binary outcome. [^8] They use this finding to motivate a careful study of the implications of the conventional definition of identification for Bayesian updating and various frequentist concepts, rather than exploring different definitions of identification as we do. In a randomized experiment, that parameter is generally partially identified in infinite population theory (Makarov 1982, Manski 1997, Fan and Park 2010), and yet the conventional definition 2 implies that it is point identified. As they note, in the binary treatment and binary instrument setting of Imbens and Angrist (1994), this implies that the proportion of compliers is point identified even without a no-defiers assumption, another inconsistency with infinite population theory. Our result goes further to show that the compliance type (i.e., complier, defier, always taker, or never taker) is point identified for every unit, according to the conventional definition of identification (see appendix LABEL:sec:conventionalAppendix), a further inconsistency with infinite population theory.
Discussion: Averaging over Units or over Counterfactual Worlds
The different conclusions implied by different definitions of identification are a direct consequence of how “data” is defined. The conventional definition 2 uses data that is not even in principle available to real-world researchers. That follows since means like $\mathbb{E}_{\text{design}}$ average over hypothetical, counterfactual worlds that are mutually incompatible with each other—such as a world where unit $i$ is treated and a world where unit $i$ is not treated. By allowing identification to depend on information from both such worlds, this definition allows one to conclude that both $Y_{i}(1)$ and $Y_{i}(0)$ are point identified. This conclusion suggests that there is, in fact, no fundamental problem of causal inference (Holland 1986).
In stark contrast, both our definition 1 and the standard definition of infinite population identification do not assume one can observe “data” from such mutually incompatible worlds. This holds for our definition 1 since the identified set only depends on $\mathbf{P}^{\text{data}}$, the matrix of data that is actually observed by real empirical researchers. The standard infinite population definition is also not based on data from incompatible worlds, and hence does not lead to results like proposition 2. Specifically, consider a joint distribution of random variables $(Y,X)$, representing the population “data” of realized outcomes and treatments. As explained in Heckman and Vytlacil (2007, section 2.2), for example, each realization $(Y(\omega),X(\omega))$ from this distribution represents a single unit $\omega$ in some set of units $\Omega$ (they assume $\Omega=[0,1]$) who is either treated ($X(\omega)=1)$ or not ($X(\omega)=0$). Consequently, even though the population is infinite, each unit is only treated once. The “data” available to researchers comes from a single assignment of each of the infinitely many units to be treated or not. Concretely, infinite population expectations like $\mathbb{E}\left[\frac{YX}{\mathbb{P}(X=1\mid W)}\right]$, as used in a selection-on-observables analysis (e.g., Imbens and Rubin 2015, section 12.4), do not involve averaging worlds where some units are treated with other worlds where those same units are not treated. Rather, such averages are over different units in the infinite population instead of over different counterfactual worlds involving different treatment assignments, as in $\mathbb{E}_{\text{design}}$.
In summary, these issues with the conventional definition motivate our development of a different theory of identification, based on definition 1.
@@ 3 Design-Based Sensitivity Analysis
Thus far we have defined and studied identification in finite populations. We also derived the identified set for ATE, $\Theta_{I}(K)$, under the $K$-approximate mean balance assumption. In this section we show how to use this identified set, combined with a known design distribution of treatment, to quantify uncertainty about ATE.
@@ 3.2 Interpreting $\Theta_{I}(K(\alpha))$
Suppose A1 and A3 hold. Then for any realization $\mathbf{X}$, $\underline{p}$ is monotonic, $\underline{p}$ is right continuous, $\lim_{K\rightarrow 0}\;\underline{p}(K)=0$, and $\lim_{K\rightarrow\infty}\;\underline{p}(K)=1$.
$K(\alpha)$ is therefore the smallest value of $K$ such that $K^{\text{true,max}}\leq K$ occurs with probability at least $1-\alpha$, according to the “robust” prior $\underline{p}$, which implies the same would be true for any other prior consistent with our baseline assumptions about the potential outcomes. So $0,K(\alpha)]$ is a Bayesian credible set for $K^{\text{true,max}}$ with respect to $\underline{p}$; it is a set which contains $K^{\text{true,max}}$ with probability about $1-\alpha$ (the true probability is slightly higher than $1-\alpha$ since $\underline{p}$ is not continuous). And, again, this would also be true for any other prior consistent with our baseline assumptions. Consequently, $\Theta_{I}(K(\alpha))$ can be interpreted as a robust $(1-\alpha)$-credible set for the identified set for ATE. Since the identified set must contain the true ATE when the model is not falsified, this implies that $\Theta_{I}(K(\alpha))$ is also a robust $(1-\alpha)$-credible set for ATE itself.
@@
Keep in mind that $\mathbf{X}^{\text{new}}$ is the only random vector here. Theorem 3 shows that $\Theta_{I}(K(\alpha))$ is a $100(1-\alpha)$% design-based confidence set. That is, across repeated random assignments of treatment, it will contain the true parameter with probability at least $1-\alpha$. In particular, this result holds for any fixed, finite population size $N$ that satisfies A1.
@@ 3.3 The Value of Randomization
    $\sup_{\theta\in\Theta_{I}(K(\alpha))}|\theta-\text{ATE}_{N}|\xrightarrow{p}0\qquad\text{as $N\rightarrow\infty$.}$
@@ 3.4 Measuring the Strength of Evidence: $\underline{p}(K^{\text{bp}})$ as an Alternative to $p$-values
$p$-values are commonly interpreted as quantitative measures of the evidence against the null hypothesis, with small values interpreted as “strong evidence” against the null and larger values interpreted as weaker evidence against it. This is a controversial interpretation (e.g., see the entire 2019 special issue of The American Statistician on “a world beyond p<0.05p<0.05”). In light of this debate about $p$-values, our measure $\underline{p}(K^{\text{bp}})$ can be viewed as an alternative summary statistic that has a well justified interpretation as a quantitative measure of how much evidence the data provides for a specific conclusion. In particular, because our analysis has a Bayesian interpretation, it allows us to make certain probabilistic statements about the true parameter itself. For example, when the naive difference in means estimate is positive, $\underline{p}(K^{\text{bp}})$ is a lower bound on the probability that the true ATE is nonnegative. Here “probability” refers to our randomization-based prior distribution.
@@ 4 Numerical Illustration
Next we use simulated data to illustrate the design-based sensitivity analysis of section 3. We generated a nested sequence of five populations, with $N\in\{10,20,40,100,400\}$. These populations have $y_{\text{min}},y_{\text{max}}]=[0,1]$ and heterogeneous treatment effects such that $\text{ATE}_{N}$ converges to 0.25 as $N$ gets large. We assigned treatment via A3 such that $N_{1}=N/2$ for all $N$ while ensuring the populations are nested. In this illustration the assignment happens only once. Appendix LABEL:sec:numericalAppendix gives further details on how we generated the data.
First we demonstrate the convergence results of theorem 4. Its first part shows that $\underline{p}(K^{\text{bp}})\xrightarrow{p}1$ as $N\rightarrow\infty$. In the proof, we showed that $\underline{p}(\cdot)$ converges to a step function at zero. Figure 2(a) demonstrates this convergence, which implies convergence of $\underline{p}(K^{\text{bp}})$ to 1 so long as the limiting breakdown point is nonzero. The second part of theorem 4 shows that, for any $\alpha\in(0,1)$, the distance between the set $\Theta_{I}(K(\alpha))$ and the ATE value obtained under point identification ($K=0$) converges to zero as $N\rightarrow\infty$. Figure 2(b) plots $\Theta_{I}(K(\alpha))$ as a function of $1-\alpha$, for the five different values of $N$. The bounds are centered at the naive difference-in-means estimand $\overline{Y}_{1}-\overline{Y}_{0}$, which varies with $N$ but converges to 0.25 by construction. Again, consistent with the theory, we see that for any $\alpha\in(0,1)$ our bounds shrink as $N$ gets larger. This plot also shows the first part of theorem 4, since $\underline{p}(K^{\text{bp}})$ is the point at which the bounds intersect the horizontal axis at zero. We see that this point converges to 1 as $N$ gets large. Indeed, for $N=100$ our bounds are strictly positive for almost all values of $1-\alpha$.